OpenStudy (anonymous):

Please help. Identify the slope and the y-intercept of the equation. . y = -2/3 x – 2

OpenStudy (anonymous):

Slope: -2/3 Y-Intercept: -2

OpenStudy (anonymous):

A linear function has the general form y=mx+c Where m (the co-efficient of x) is the slope/grdient anf the constant c is the y-intercept

OpenStudy (anonymous):

-2/3 = slope or m, y = -2... (-2,0) is the coordinate of y intercept or y = -2 or b...

OpenStudy (anonymous):

One of the general forms of linear equation is y=mx+c ,where m=slope and c=y-inercept let we compare your given equation ( y = -2/3 x – 2) to the general one. -2/3 x -2 +mx+c so m=slope=-2/3 and c=y-intercept=-2

OpenStudy (anonymous):

its quite easy y intercet is -2 slope is -2/3

OpenStudy (anonymous):

put x=0 to find the y-intercept so y inttercept is -2 ........ while if u want to find the x-intercept put y=0 in the eq.